# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ A.12 Client-Wise Accuracy
We first compare client-wise performance in the homogeneous PFL-Dynamic setups on DRAKE and HFLB, and summarize the results in Tab.15 and Tab.16, respectively. As shown in the tables, personalization performance is improved for all clients compared to SFT (i.e., supervised fine-tuning on local data) except Client 8 in Tab. 16. We highlight this result because SFT is a strong baseline for personalization (ghari2024personalized, ), as we also show in Tab. 13, where it outperforms all baselines except for FedMosaic. Moreover, supervised fine-tuning may be sufficient for personalization in some tasks (wozniak2024personalized, ; mosbach2021on, ), as seen with Client 6 in Tab. 15, where there are only marginal improvements compared to SFT. However, for more challenging tasks, properly leveraging knowledge from other clients significantly improves personalization, e.g., FedMosaic shows a gain of 8.2% in $A_{last}$ for Client 9 and 6.4% for Client 7, as well as a 7.1% improvement in $A_{\text{AUC}}$ for Client 5. In summary, even though there is only a 1–3% gain in the average performance across clients, this is because in clients where SFT is sufficient for personalization, the gain from PFL seems smaller. Consistent with the PFL-Dynamic setup, FedMosaic outperforms SFT in the PFL-Static scenario, as shown in Tab. 17.
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
Method $A_{last}$ / $A_{\text{AUC}}$
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
A.13 Effect of Low Rank $r$ in PQ-LoRA
@@
We study how varying the rank $r$ in PQ-LoRA affects performance, and summarize the results in Fig. 10. As shown in the figure, while both excessively high and low values of $r$ lead to degraded performance, a broad range of intermediate values shows stable performance. While a larger $r$ increases shareable capacity ($\mathbb{R}^{r\times r}$), it may cause overfitting (yang2024lora, ; cho2024heterogeneous, ). In contrast, a smaller $r$ helps mitigate overfitting but may introduce capacity limitations, resulting in suboptimal performance (he2022towards, ). By balancing the trade-off, we select an adequate low rank $r$.
@@
Figure 10: Accuracy on different low rank $r$ values in PQ-LoRA in PFL-Dynamic setup on DRAKE.
A.14 Ablation Study of SIMA
@@
Our proposed SIMA measures task similarity among clients using client-wise gradients to construct a similarity-aware customized global model for each client. Specifically, for the $i$-th client, we maintain a decayed gradient $\hat{g}_{i}$, updated via exponential moving average (EMA) from the current gradient $g_{i}$, which is computed using the current batch and a small, frozen pre-trained model. Note that we use the last-layer gradient of the frozen model, not the training model, as mentioned in Sec. 4.1.
@@
We ablate SIMA by comparing four aggregation variants: (1) Equal-weight aggregation, which ignores task similarity and aggregates local models uniformly; (2) Training model gradients, using current batch gradients from the trainable model; (3) Frozen model gradients, using current batch gradients from the frozen pre-trained model; and (4) Decayed frozen gradient (i.e., SIMA), which maintains an EMA of gradients from the frozen pre-trained model to capture model knowledge shifts. We summarize the results in Tab. 20.
@@
As shown in the table, model aggregation considering task similarity measured by client-wise gradients generally improves the ‘Self’ accuracy. This demonstrates the loss of information in Equal-weight aggregation due to parameter interference. While using gradients from the training model (‘Training model gradient’) shows improved performance compared to Equal-weight aggregation, it still suffers from data heterogeneity, as gradient similarities from models trained on heterogeneous data may not capture actual similarity (tang2020similarity, ; evans2024data, ). Using a frozen model’s gradients from the current batch (‘Frozen model gradients’) can address this limitation, but cannot reflect model knowledge shifts under distribution shifts, as mentioned in Sec. 4.1. In contrast, SIMA effectively captures task similarity under data heterogeneity with distribution shifts, thus outperforming other aggregation strategies.
Method Self
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Equal-weight aggregation 66.99$\pm$0.77 58.39$\pm$0.38
Training model gradients 67.16$\pm$0.74 59.38$\pm$0.53
Frozen model gradients 67.24$\pm$1.27 59.45$\pm$0.74
Decayed frozen gradients (SIMA) 67.86$\pm$0.51 59.83$\pm$0.16
@@
Table 20: Ablation of components in SIMA in Heterogeneous PFL-dynamic setup on DRAKE. ‘Equal-weight aggregation’ refers to aggregating local models with equal weight, ‘Training model gradients’ refers to using gradients from the current batch computed with the training model, ‘Frozen model gradients’ refers to using gradients from the current batch computed with the frozen pre-trained model, and ‘Decayed frozen gradients’ refer to EMA-estimated gradients from the frozen model (i.e., SIMA).
A.15 Different Weight Alignment Methods in PQ-LoRA
@@
To ensure shared initialization across heterogeneous architectures, we align $A$ matrices from heterogeneous architectures (i.e., $A_{i}\in\mathbb{R}^{r\times d_{i}}$, $A_{j}\in\mathbb{R}^{r\times d_{j}}$) using MSE loss, and $B$ matrices (i.e., $B_{i}\in\mathbb{R}^{d^{\prime}_{i}\times r}$, $B_{j}\in\mathbb{R}^{d^{\prime}_{j}\times r}$) using canonical correlation analysis (CCA), as detailed in Sec.4.2.2. We ablate the effects of aligning $A$ and $B$ matrices, and summarize the results in Tab. 21. Note that the results in the table report the average performance across all clients, which can make the overall average improvements appear small, as mentioned in Sec. 6.2.
@@
As shown, alignment improves both personalization (‘Self’) and generalization (‘Others’) performance. This occurs because aligning both $A$ and $B$ in PQ-LoRA ensures heterogeneous models share initialization and follow consistent optimization paths, allowing aggregation without weight interference (wortsman2022model, ; wortsman2022robust, ; yadav2023ties, ). In contrast, misalignment in either matrix introduces initialization discrepancies between models, leading to weight interference during aggregation and degrading performance (jordanr2023epair, ; neyshabur2020being, ; stoica2025model, ).
@@
A B $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Random Random 67.95$\pm$1.71 63.50$\pm$0.66 47.71$\pm$0.08 47.46$\pm$0.05
Random Aligned (CCA) 68.22$\pm$1.84 63.59$\pm$0.62 47.80$\pm$0.06 47.56$\pm$0.08
Aligned (MSE) Random 68.29$\pm$1.57 63.58$\pm$0.60 47.72$\pm$0.14 47.41$\pm$0.04
Aligned (MSE) Aligned (CCA) 68.73$\pm$1.92 63.82$\pm$0.59 48.04$\pm$0.04 47.57$\pm$0.08

### v3

@@ A.16 Client-Wise Accuracy
We first compare client-wise performance in the homogeneous PFL-Dynamic setups on DRAKE and HFLB, and summarize the results in Tab.21 and Tab.22, respectively. As shown in the tables, personalization performance is improved for all clients compared to SFT (i.e., supervised fine-tuning on local data) except Client 8 in Tab. 22. We highlight this result because SFT is a strong baseline for personalization (Ghari & Shen, 2024), as we also show in Tab. 17, where it outperforms all baselines except for FedMosaic. Moreover, supervised fine-tuning may be sufficient for personalization in some tasks (Woźniak et al., 2024; Mosbach et al., 2021), as seen with Client 6 in Tab. 21, where there are only marginal improvements compared to SFT. However, for more challenging tasks, properly leveraging knowledge from other clients significantly improves personalization, e.g., FedMosaic shows a gain of 8.2% in $A_{last}$ for Client 9 and 6.4% for Client 7, as well as a 7.1% improvement in $A_{\text{AUC}}$ for Client 5. In summary, even though there is only a 1–3% gain in the average performance across clients, this is because in clients where SFT is sufficient for personalization, the gain from PFL seems smaller. Consistent with the PFL-Dynamic setup, FedMosaic outperforms SFT in the PFL-Static scenario, as shown in Tab. 23.
$_{last} / AA$$_{\text{AUC}}$
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
$_{last} / AA$$_{\text{AUC}}$
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
$_{last} / AA$$_{\text{AUC}}$
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
@@
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9
@@
A.17 Details of Public Data $D_{p}$
For the public data $D_{p}$, we use a subset of the MLLM’s pretraining data, as briefly mentioned in Sec. 4.2.2. Specifically, since our experiments are based on LLaVA, we adopt its instruction tuning dataset, LLaVA-Instruct-158K (Liu et al., 2023b), as $D_{p}$. For the alignment details, we randomly sample 5,000 examples from $D_{p}$ and align PQ-LoRA for 1 epoch using a batch size of 4, a learning rate of $5\times 10^{-5}$, and the Adam optimizer with cosine scheduler. To further assess the effect of $D_{p}$, we explore the effect of $D_{p}$’s size and its distributional alignment with the client data.
We first assess the effect of public data size in Tab 26. As shown, using too little data results in insufficient alignment of PQ-LoRA, leading to degraded performance. Interestingly, using too much public data also causes slight performance drops, likely due to overfitting to the $D_{p}$ distribution, which hinders generalizable alignment to clients’ distributions. Considering this trade-off, we selected a moderate dataset size (i.e., 5,000 samples).
Next, we assess robustness under domain mismatch between $D_{p}$ and the clients’ distribution. Specifically, we use ChatbotIT (Zheng et al., 2023) (text-only NLP benchmark) and Visual Storytelling (Li et al., 2024d) (multi-sentence outputs) as $D_{p}$, both differing from the clients’ multi-modal, short-answer tasks. To ensure a fair comparison, we fix the size of $D_{p}$ to 5,000 samples. As shown in Tab. 27, using Visual Storytelling shows comparable performance with LLaVA-Instruct-158K. Similarly, using text-only NLP benchmark shows performance comparable to (and even surpassing) LLaVA-Instruct-158K, demonstrating resilience of FedMosaic to distribution misalignment between $D_{p}$ and clients’ data.
Self Others
Size of $D_{p}$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 26: Effect of public data $D_{p}$ size. FedMosaic performs best on moderate-sized datasets (i.e., 1,250-5,000)
Self Others
Dataset $D_{p}$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 27: Effect of domain gap between $D_{p}$ and clients’ data. FedMosaic shows consistent performance across multiple benchmarks, even on text-only NLP benchmarks (i.e., ChatbotIT).
A.18 Hyperparameter Analysis
Since dataset-specific hyperparameter search is undesirable under distribution shifts (i.e., Dynamic setup), where future data are unknown, we adopt a single set of hyperparameters across all benchmarks and setups, determined from the DRAKE-Dynamic setup.
Effect of Communication Rounds $R$ and Local Steps.
We analyze the trade-off by varying the number of communication rounds and local steps under a fixed training budget. As shown in Tab. 28, more rounds (i.e., fewer local steps per round) degrade personalization (‘Self’). We attribute this to insufficient local adaptation, which weakens the personalized models and reduces the quality of shared knowledge during communication, ultimately limiting improvements in generalizability. Conversely, significantly reducing rounds (i.e., increasing local steps) improves personalization but harms generalization (‘Others’) due to limited inter-client knowledge sharing, reducing generalizability. To balance this trade-off, we adopt a moderate setting with 20 rounds and 100 local steps per round.
Self Others
Total Rounds Local Steps per Round $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 28: Effect of communication rounds and local steps on DRAKE under Fixed Total Training Cost (i.e., Total rounds $\times$ Local steps per round).
Effect of Low Rank $r$ in PQ-LoRA
@@ A.18 Hyperparameter Analysis
We study how varying the rank $r$ in PQ-LoRA affects performance, and summarize the results in Fig. 13. As shown in the figure, while both excessively high and low values of $r$ lead to degraded performance, a broad range of intermediate values shows stable performance. While a larger $r$ increases shareable capacity ($\mathbb{R}^{r\times r}$), it may cause overfitting (Lin et al., 2024; Cho et al., 2024). In contrast, a smaller $r$ helps mitigate overfitting but may introduce capacity limitations, resulting in suboptimal performance (He et al., 2022). By balancing the trade-off, we select an adequate low rank $r=128$.
@@
Figure 13: Accuracy on different low rank $r$ values in PQ-LoRA on DRAKE-Dynamic setup. Extremely low or high ranks degrade performance, but a wide intermediate range maintains stable accuracy.
Effect of $N_{B}$ in PQ-LoRA
We study the effect of $N_{B}$, the number of layers employing PQ-LoRA, and summarize the results in Tab.29. In a model $W$ with $|W|$ layers, increasing $N_{B}$ reduces the layers using conventional LoRA to $|W|-N_{B}$. Since PQ-LoRA has fewer trainable parameters (i.e., $r^{2}+r$) than conventional LoRA (i.e., $r\times(d_{I}+d_{O})$), increasing $N_{B}$ decreases the total number of learnable parameters in $W$, as $A\in\mathbb{R}^{r\times d_{I}}$ and $B\in\mathbb{R}^{d_{O}\times r}$ are frozen, with only $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$ being trainable. Despite the reduced number of parameters, as shown in the SFT performance in the table, performance remains stable even as $N_{B}$ increases. We attribute this to orthogonal and frozen $A$ and $B$, which maximize and preserve the expressiveness of PQ-LoRA (Theorem1), despite the lower trainable parameter count.
Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Effect of Orthogonality-regularization Scale $\lambda$ in PQ-LoRA
We study how varying the orthogonality-regularization scale $\lambda$ in PQ-LoRA alignment process (Eq. 20) affects performance, and summarize the results in Fig. 14. We observe that the ‘Self’ performance does not fluctuate on larger $\lambda$, but reduces for small $\lambda$. This indicates that insufficient weighting of the orthogonality-regularization term weakens the enforced orthogonality among $A$ matrices from heterogeneous models, thereby reducing the representational capacity (i.e., Span) of PQ-LoRA, as mentioned in Theorem 1.
Figure 14: Accuracy on different L2 regularization scale $\lambda$ values in PQ-LoRA on DRAKE-Dynamic setup. Extremely low $\lambda$ degrades performance, but a wide intermediate range maintains stable accuracy.
Effect of Temperature $\tau$ in RELA
We analyze the effect of the softmax temperature $\tau$ in RELA (Eq. 4), which controls the aggregation sharpness, i.e., lower $\tau$ focuses aggregation on similar clients, while higher $\tau$ aggregates weights across more diverse ones. Interestingly, as shown in Fig. 15, neither decreasing nor increasing the temperature $\tau$ consistently improves personalization or generalization. Instead, we observe two trends: (i) Extremely low $\tau$ degrades both personalization and generalization (ii) Increasing $\tau$ too much reduces personalization, while generalization gradually stabilizes.
This is because low $\tau$ limits the diversity of aggregation, restricting knowledge sharing even among moderately relevant clients. In contrast, high $\tau$ encourages sharing across dissimilar tasks, which can reduce personalizability due to the inclusion of unrelated knowledge in the global model. However, increasing $\tau$ does not continuously improve generalization; it converges after a certain point. We believe this is because, although a higher $\tau$ promotes aggregation of more diverse knowledge, it also increases parameter interference (e.g., sign conflicts) when combining models trained on distinct tasks (Yeh et al., 2023; Ding et al., 2024). Consequently, we adopt a moderate temperature of $\tau=0.5$ for all experiments.
Figure 15: Accuracy under different temperatures $\tau$ on DRAKE-Dynamic setup. Extremely low or high temperatures degrade performance, but a wide intermediate range maintains stable accuracy.
Effect of Gradient sampling ratio $N_{s}$ in RELA
We analyze the effect of gradient sampling ratio $N_{s}$ and summarize the results in Fig. 16. As shown, sampling only a very small number of dimensions from the decayed client-wise gradient $\hat{g}_{i}$ to construct the sanitized gradient $\tilde{g}_{i}$ significantly degrades performance, as it undermines the representativeness and informativeness of the compressed gradient vector (Li et al., 2024b). However, using up to $N_{s}=40\%$ of the gradient dimensions maintains comparable performance while reducing both communication costs and privacy risks from gradient inversion. Accordingly, we consistently set $N_{s}=40\%$ for RELA in all experiments.
Figure 16: Accuracy under different gradient sampling ratio $N_{s}$ on DRAKE-Dynamic setup. Extremely low gradient sampling ratio degrades performance, but a wide range maintains stable accuracy.
Effect of decaying EMA ratio $\alpha$ in RELA
We analyze the effect of gradient deacying EMA ratio $\alpha$ and summarize the results in Fig. 17. Too small $\alpha$ ignores current task information, while too large $\alpha$ ignores previously learned tasks. Balancing the trade-off, we consistently set $\alpha=0.5$ for RELA in all experiments.
Figure 17: Accuracy under different gradient decaying EMA ratio $\alpha$ on DRAKE-Dynamic setup. Balance between current and old ($\alpha=0.5$) shows the best result.
Effect of noise scale $\mu$ in RELA sanitized gradient
We also study the noise scale $\mu$ in sanitized gradient $\tilde{g}$ (Eq. 3) in RELA and visualize the results in Fig. 18. We clearly see the decreasing trend in the results when we increase the noise scale. While stronger random noise enhances privacy by better concealing sensitive information, excessive noise severely distorts the gradient signal, undermining the reliable estimation of task relevance based on gradient similarity.
Figure 18: Accuracy under different noise scale $\mu$ on DRAKE-Dynamic setup. Strong noise distorts the gradient signal, undermining the reliable estimation of task relevance based on gradient similarity.
A.19 Ablation Study of RELA
@@
Our proposed RELA measures task similarity among clients using client-wise gradients to construct a similarity-aware customized global model for each client. Specifically, for the $i$-th client, we maintain a decayed gradient $\hat{g}_{i}$, updated via exponential moving average (EMA) from the current gradient $g_{i}$, which is computed using the current batch and a small, frozen pre-trained model. Note that we use the last-layer gradient of the frozen model, not the training model, as mentioned in Sec. 4.1.
@@
We ablate RELA by comparing four aggregation variants: (1) Equal-weight aggregation, which ignores task similarity and aggregates local models uniformly; (2) Training model gradients, using current batch gradients from the trainable model; (3) Frozen model gradients, using current batch gradients from the frozen pre-trained model; and (4) Decayed frozen gradient (i.e., RELA), which maintains an EMA of gradients from the frozen pre-trained model to capture model knowledge shifts. We summarize the results in Tab. 30.
@@
As shown in the table, model aggregation considering task similarity measured by client-wise gradients generally improves the ‘Self’ accuracy. This demonstrates the loss of information in Equal-weight aggregation due to parameter interference. While using gradients from the training model (‘Training model gradient’) shows improved performance compared to Equal-weight aggregation, it still suffers from data heterogeneity, as gradient similarities from models trained on heterogeneous data may not capture actual similarity (Tang et al., 2020; Evans et al., 2024). Using a frozen model’s gradients from the current batch (‘Frozen model gradients’) can address this limitation, but cannot reflect model knowledge shifts under distribution shifts, as mentioned in Sec. 4.1. In contrast, RELA effectively captures task similarity under data heterogeneity with distribution shifts, thus outperforming other aggregation strategies.
Self
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
@@
Table 30: Ablation of components in RELA in Heterogeneous PFL-dynamic setup on DRAKE. ‘Equal-weight aggregation’ refers to aggregating local models with equal weight, ‘Training model gradients’ refers to using gradients from the current batch computed with the training model, ‘Frozen model gradients’ refers to using gradients from the current batch computed with the frozen pre-trained model, and ‘Decayed frozen gradients’ refer to EMA-estimated gradients from the frozen model (i.e., RELA).
A.20 RELAwith Lighter $W_{s}$
To extract per-client gradients for RELA, we employ the smallest MLLM among all clients’ models (e.g., LLaVA-1.5-1B). To further reduce the computational overhead of RELA, we can utilize an even smaller model (e.g., LLaVA-Qwen-0.5B) or apply lower-bit quantization (e.g., 4-bit and 8-bit), as shown in Tab. 31. The results show that FedMosaic maintains consistent performance even with smaller models and lower-bit quantization, demonstrating that computational costs can be further reduced without sacrificing performance.
Self Others
$W_{s}$
Relative computational costs $\downarrow$
$_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 31: Effect of $W_{s}$’s model size and quantization on DRAKE-Dynamic. Relative computational costs denote the ratio of computation compared to LLaVA-1.5-1B (16-bit). Comparable performance is maintained with both a smaller model (i.e., 0.5B) and quantized models (i.e., 4-bit, 8-bit).
A.21 Different Weight Alignment Methods in PQ-LoRA
@@
To ensure shared initialization across heterogeneous architectures, we align $A$ matrices from heterogeneous architectures (i.e., $A_{i}\in\mathbb{R}^{r\times d_{i}}$, $A_{j}\in\mathbb{R}^{r\times d_{j}}$) using L2 loss, and $B$ matrices (i.e., $B_{i}\in\mathbb{R}^{d^{\prime}_{i}\times r}$, $B_{j}\in\mathbb{R}^{d^{\prime}_{j}\times r}$) using canonical correlation analysis (CCA), as detailed in Sec.4.2.2. We ablate the effects of aligning $A$ and $B$ matrices, and summarize the results in Tab. 32. Note that the results in the table report the average performance across all clients, which can make the overall average improvements appear small, as mentioned in Sec. 6.2.
@@
As shown, alignment improves both personalization (‘Self’) and generalization (‘Others’) performance. This occurs because aligning both $A$ and $B$ in PQ-LoRA ensures heterogeneous models share initialization and follow consistent optimization paths, allowing aggregation without weight interference (Wortsman et al., 2022a; b; Yadav et al., 2023). In contrast, misalignment in either matrix introduces initialization discrepancies between models, leading to weight interference during aggregation and degrading performance (Jordan et al., 2023; Neyshabur et al., 2020; Stoica et al., 2025).
@@
A B $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]